OpenStudy (kittiwitti1):

\[\text{If }a\ne b\text{ and }\frac{1}{x}+\frac{1}{a}=\frac{1}{b},\text{ then }x=\]

hartnn (hartnn):

what have u tried?

OpenStudy (kainui):

@dan815

OpenStudy (kittiwitti1):

I don't even know what I did, I'm just confused

OpenStudy (dan815):

haha how did this happen

OpenStudy (kittiwitti1):

Attempt:\[\frac{ab}{z}+\frac{xb}{a}=\frac{ax}{a}\]\[b+\frac{xb}{a}=x\]

OpenStudy (dan815):